# TaeBench: Improving Quality of Toxic Adversarial Examples

### v1

@@ 1 Introduction
This careful filtering process helps us curate a high-quality dataset of more than 260k TAE examples. We name it as TaeBench (Toxic Adversarial Example Bench). There exist many potential use cases of TaeBench. In our experiments, first, we showcase one main use case as transfer attack based benchmarking. We attack SOTA toxicity content moderation models and API services using TaeBench and show they are indeed vulnerable to TaeBench with attack success rates (ASR) up to $77\%$. We then empirically show how vanilla adversarial training using TaeBench can help increase the robustness of a toxicity detector even against unseen attacks by decreasing the ASR from $75\%$ to lower than $15\%$. To the best of our knowledge, our research is the first in the literature to provide quantitative analyses and quality improvement towards building more robust toxicity guardrails.
@@ 2 Toxic Adversarial Examples (TAE) and Attack Recipes
This paper focuses on the TAE proposed by Bespalov et al. (2023) Bespalov et al. (2023). The main motivation of TAE attacks is that a major goal of real-world toxicity detection is to identify and remove toxic language. Adversarial attackers against toxicity detectors will focus on designing samples that are toxic in nature but can fool a target detector into making benign prediction (aka TAE). TAE attacks search for an adversarial example ${\mathbf{x}}^{\prime}$ from a seed input ${\mathbf{x}}$ by satisfying a targeted goal function as follows: ${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime})=\{{\mathcal{F}}({\mathbf{x}}^{\prime})=b;{\mathcal{F}}({\mathbf{x}})\neq b\}$ (Eq-1). Here $b$ denotes the "0:benign" class. ${\mathcal{F}}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is a given target toxicity text classifier.
Adversarial attack methods design search strategies to transform a seed ${\mathbf{x}}$ to ${\mathbf{x}}^{\prime}$ via transformation, so that ${\mathbf{x}}^{\prime}$ fools ${\mathcal{F}}$ by achieving the fooling goal ${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime})$, and at the same time fulfilling a set of constraints. Therefore literature has split each text adversarial attack into four components: (1) goal function, (2) transformation, (3) search strategy, and (4) constraints between seed and its adversarial examples Morris et al. (2020a). This modular design allows pairing the TAE goal function (Eq-1) with popular choices of other three components from the literature to obtain a large set of TAE attack recipes.
@@ 2.1 Running $>20$ SOTA Recipes for a Large Unfiltered TAE Pool
[…]
Our first step is to select a total of 25 TAE attack recipes to generate a large pool of raw TAE samples (see Section 4 for details of two seed datasets and three proxy toxicity detection models). Specifically, we use 20 variants of attack recipes proposed in ToxicTrap Bespalov et al. (2023) that combine different transformation, constraint, and search strategy components. In addition to these ToxicTrap attack recipes, we select 5 algorithms from literature: DeepWordBug Gao et al. (2018), TextBugger Li et al. (2019), A2T Yoo and Qi (2021), PWWS Ren et al. (2019), and TextFooler Jin et al. (2019). These algorithms were proposed to attack general language classifiers. We adapt these five attacks by replacing their goal functions with Eq-1. These 25 attack recipes cover a wide range of popular transformations, constraints, and search methods (details in Table A1).
@@ 3.1 LLM Judge and Small Models based Automated Quality Controls
Removing text of low naturalness. Next we remove samples with low text naturalness using an English acceptability classifier (Proskurina et al., 2023). This classifier is fine-tuned from Huggingface TDA-BERT using a $3k$ labeled data we collect through human annotation. The human annotation guidelines on what defines "text naturalness" are in Section 3.2. We fine-tune the model with $2,370$ labeled texts, and evaluate it with $593$ held-out texts, following training setup in Section C.3. Table A3 shows that the F1 score ($88.9\%$) of fine-tuned TDA-BERT improves 18% compared to F1 ($70.5\%$) from pretrained TDA-BERT.
LLM judge for Removing non-toxic invalid TAE samples. Now we design model-based automated strategy to keep only those TAE samples that are semantically toxic. We propose an ensemble approach for toxicity label filtering by combining : (1) in-context learning (ICL) prompted Mistral (Mistral-7B-Instruct-v0.1) Jiang et al. (2023) and (2) a fine-tuned toxigen-RoBERTa classifier Hartvigsen et al. (2022) (via "AND"). For (1), Mistral ICL, we run a series of experiments to select the best ICL prompt formatting according to He et al. (2024) and build 5-shot ICL prompting by selecting demonstrations from our TAE dataset (see the prompt in Table A5). The accuracy of best Mistral ICL prompting is 76%. For (2), we fine-tune Toxigen-Roberta with $3.2k$ human annotated data (see annotation guideline in Section C.2 and training set up in Section C.3) and achieve a F1 score of 94% (Table A4).
@@ 3.2 Human Evaluation to Annotate TAE on Toxicity and Naturalness
We use human annotators to curate the toxicity and text naturalness of subsets of generated TAE examples. Three human annotators are asked to review the toxicity and three annotators are asked to annotate the text naturalness. The final label is assigned by unanimous vote, where a fourth adjudicator resolves any disagreements. (1) Toxicity is defined as "issues that are offensive or detrimental, including hate speech, harassment, graphic violence, child exploitation, sexually explicit material, threats, propaganda, and other content that may cause psychological distress or promote harmful behaviors." (2) Text naturalness is defined as "text that could be plausibly written by a human even if it includes ’internet language’ that is outside ’school grammar’"
@@
Dataset Seeding Source Train Test Jigsaw Jig (2018) - 1.48MM 185k Off-Tweet Davidson et al. (2017a) - 20k 2.5k Raw TAEs Jigsaw 529,880 271,805 OffensiveTweet 57,639 77,418 TaeBench Jigsaw 197,734 38,539 OffensiveTweet 12,857 15,989 TaeBench+ Jigsaw 199,244 40,114 OffensiveTweet 13,837 16,115
@@ 4.1 TAE Generation with Proxy Models and Seeding Datasets
Running TAE attacks needs a set of text inputs that are toxic as seeds (denoted as ${\mathbf{x}}$ in Eq-1 of Section 2.1). We use the following two datasets as seeds for our TAE attacks.
@@
A dataset derived from the Wikipedia Talk Page dataset Jig (2018). Wikipedia Talk Page allows users to comment, and the comments are labeled with toxicity levels. Comments that are not assigned any of the six toxicity labels are categorized as "non toxic". We can use this data for both binary and multi-label toxicity detection tasks.
@@
Davidson et al. (2017) Davidson et al. (2017b) used a crowd-sourced hate speech lexicon from Hatebase.org to collect tweets containing hate speech keywords. Each sample is labeled as one of three classes: those containing hate speech, those containing only offensive language, and those containing neither. This data is for multi-class toxicity detection.
@@
One important property of adversarial attacks is the ability of the attack to transfer from the model used in its development to attacking other independent models. Transferability occurs because deep learning models often learn similar decision boundaries and features. Therefore, perturbations and noise patterns that fool one model are likely to also fool other models trained on the same or similar datasets. Motivated by adversarial transferability, we build three local text toxicity models as target proxies and run 25 different TAE attack recipes (see Section 2.1) against them to generate a large-scale pool of unfiltered TAE dataset (940k samples in total). Details of these proxy models are in Table A2 and Section C.4.
@@ 4.2 TaeBench: a Large Set of Quality Controlled TAE Samples
In Table 2, we pass a total of 936,742 raw TAEs through the proposed quality filtering pipeline. We are able to select 264,672 examples (28.30% as of the original examples) as the filtered set, and we call it TaeBench.
To validate the quality of each step of filtering, we conduct human-in-the-loop annotations by randomly sampling 200 TAEs from each filtering step to evaluate the ratios of toxic and natural examples. In Table 2, human validation shows that, after filtering, the toxicity ratios are improved by 5.64% in the selected examples ($94.17\%$) compared to unfiltered examples ($88.53\%$). The text naturalness ratios are improved by 6.36%, from ($79.63\%$) in the unfiltered examples to ($85.99\%$) in the selected examples.
@@ 4.3 TaeBench+: Benign Seeds Derived Adversarial Examples
TAE are semantic-toxic samples that fool toxicity detection models into making benign predictions. Essentially they are false negative predictions (assuming "toxic" is the positive class). Related, it is also interesting to understand and search for those semantic-benign samples that fool a target model into making toxic predictions. These samples belong to false positive inputs. We call them "benign adversarial examples (BAE)" in the rest of this paper.
To search for BAE, we design its goal function as: ${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime})=\{{\mathcal{F}}({\mathbf{x}}^{\prime})\neq b;{\mathcal{F}}({\mathbf{x}})=b\}$ (Eq-2), where $b$ denotes the benign class. Starting from benign seeds (${\mathcal{F}}({\mathbf{x}})=b$), we perturb ${\mathbf{x}}$ into ${\mathbf{x}}^{\prime}$ by pushing the prediction of ${\mathbf{x}}^{\prime}$ to not be benign anymore. We can reuse the TAE attack recipes by keeping their transformation, search and constraint components intact, and replace the goal function into the above equation (Eq-2).
Empirically, we run the 25 BAE attacks following the same setup as TaeBench, obtaining 102,667 raw BAE examples (searching for BAE seems harder than searching for TAE). Table A8 shows how we conduct automated filtering following the same workflow as obtaining TaeBench. Differently, in the label-toxicity filtering step, we keep those benign-labeled BAE samples. Finally, we add the filtered BAE examples to create TaeBench+, a new variation of the TaeBench dataset. We provide the additional benefits of TaeBench+ in Section 5.3.
@@ 5.1 Benefit I: Benchmark Toxicity Detectors via Transfer Attacks
To evaluate the efficacy of the filtered TAE examples, we conduct transfer attack experiments to benchmark four SOTA toxicity classifiers: detoxify (detoxify-unbiased) (Hanu and Unitary team, 2020), Llama Guard333meta-textgeneration-llama-guard-7b (Inan et al., 2023), OpenAI Moderation API444text-moderation-007 from https://platform.openai.com/docs/guides/moderation/overview, and Nemo Guardrails (with GPT-3.5-turbo) (Rebedea et al., 2023). Using TaeBench in transfer attacks can save resources and minimize the effort needed to generate TAE examples plus with data quality guarantees. Also the transfer attack set up is indeed a (major) real-world use case of using TAE. In this black-box transfer attack setup, TAE are constructed offline (like what we have done using many existing TAE attack recipes to attack local proxy models), then get them used to attack a target victim model or service.
@@
5.2 Benefit II: Improve Toxicity Detectors w. Adversarial Training
We also showcase how a vanilla adversarial training with TaeBench can help increase the adversarial robustness of a toxicity detector even against unseen attacks. Here, adversarial training introduces the TAE adversarial data into the training of a DistilBERT or detoxify model together with the Jigsaw Binary train split. More dataset details are in Table 3.
In Table 5, we report the impacts of using TaeBench for adversarial training. We train DistilBERT/detoxify models using (a) Jigsaw-train only (No TAE); (b) Jigsaw-train + extra unfiltered TAE samples (TAE-Unfiltered); and (c) Jigsaw-train + TaeBench. We sample the unfiltered TAE data such that TAE-Unfiltered has the same size as TaeBench to have a fair comparison on model performance by removing the impact of data set size. We observe that the model trained with Jigsaw-train + TaeBench achieves significantly lower ASR (14.58% and 23.25% FNR for DistilBERT and detoxify, respectively), being more robust than the model trained with no adversarial training (74.99% and 54.28% ASR/FNR). It is also better than or close to the one trained with random sampling augmentation (16.55% and 22.92% ASR/FNR). We also observe that these three data augmentations impact the classification metrics from the Jigsaw test set very minimally (<2% impact on F1 and AUC scores in Table 5). Training setups are described in Section C.3.
@@ 5.3 Variation: Adding TaeBench+
In Table 5, when we augment the training data with TaeBench+, the model achieves the lowest ASR (FNR) of 12.66% and 22.80% on TaeBench-test for DistilBERT and detoxify, respectively. We further oversample the benign adversarial examples in TaeBench+ during data augmentation. We name it as balanced TaeBench+ and it aims to balance off the size difference between toxic and benign adversarial examples in TaeBench+. This revised augmentation helps the trained model achieve the lowest or second lowest ASR (FPR) of 53.02% and 3.97% on the (TaeBench+)-test-benign samples. ASR of running the BAE examples is essentially false positive rate (FPR). Finally, we combine the FPR and FNR into a summary metric: calculate balanced accuracy (BACC) on TaeBench+ test set. The model trained with the balanced TaeBench+ achieves the highest balanced accuracy of 65.35% and 86.55% shown in the last column of Table 5.

### v2

@@ 1 Introduction
This careful filtering process helps us curate a high-quality dataset of more than 260k TAE examples. We name it as TaeBench (Toxic Adversarial Example Bench). There exist many potential use cases of TaeBench. In our experiments, first, we showcase one main use case as transfer attack based benchmarking. We attack SOTA toxicity content moderation models and API services using TaeBench and show they are indeed vulnerable to TaeBench with attack success rates (ASR) up to $77\%$. We then empirically show how vanilla adversarial training using TaeBench can help increase the robustness of a toxicity detector even against unseen attacks by decreasing the ASR from $75\%$ to lower than $15\%$.
@@ 2 Toxic Adversarial Examples (TAE) and Attack Recipes
This paper focuses on the TAE proposed by Bespalov et al. (2023). The main motivation of TAE attacks is that a major goal of real-world toxicity detection is to identify and remove toxic language. Adversarial attackers against toxicity detectors will focus on designing samples that are toxic in nature but can fool a target detector into making benign prediction (aka TAE). TAE attacks search for an adversarial example ${\mathbf{x}}^{\prime}$ from a seed input ${\mathbf{x}}$ by satisfying a targeted goal function as follows:
${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime}):=\{{\mathcal{F}}({\mathbf{x}}^{\prime})=b;{\mathcal{F}}({\mathbf{x}})\neq b\}$ (1)
Here $b$ denotes the "0:benign" class. ${\mathcal{F}}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is a given target toxicity text classifier.
Adversarial attack methods design search strategies to transform a seed ${\mathbf{x}}$ to ${\mathbf{x}}^{\prime}$ via transformation, so that ${\mathbf{x}}^{\prime}$ fools ${\mathcal{F}}$ by achieving the fooling goal ${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime})$, and at the same time fulfilling a set of constraints. Therefore literature has split each text adversarial attack into four components: (1) goal function, (2) transformation, (3) search strategy, and (4) constraints between seed and its adversarial examples Morris et al. (2020a). This modular design allows pairing the TAE goal function (Equation (1)) with popular choices of other three components from the literature to obtain a large set of TAE attack recipes.
@@ 2.1 Running $>20$ SOTA Recipes for a Large Unfiltered TAE Pool
The research community still lacks a systematic understanding of the adversarial robustness of SOTA toxicity text detectors. Two major challenges exist: (1) running TAE attack recipes is quite time consuming; and (2) many generated TAE samples are invalid or ambiguous (see  Table 1). For instance,  Table A2 shows that the average runtime cost of running ToxicTrap Bespalov et al. (2023) attack recipes against a binary toxicity classifier from $185k$ seed samples takes ~$29.9$ hours. It takes ~$6.6$ hours to attack a multi-class toxicity detector from $2.5k$ seeds. To address this, we aim to develop a standardized, high-quality dataset of TAE examples that covers a wide range of possible attack recipes.
Our first step is to select 25 TAE attack recipes to generate a large pool of raw TAE samples (see Section 4 for seed datasets and three proxy toxicity detection models). Specifically, we use 20 variants of attack recipes proposed in ToxicTrap Bespalov et al. (2023) that combine different transformation, constraint, and search strategy components. In addition to these ToxicTrap attack recipes, we select 5 algorithms from literature: DeepWordBug Gao et al. (2018), TextBugger Li et al. (2019), A2T Yoo and Qi (2021), PWWS Ren et al. (2019), and TextFooler Jin et al. (2019). These algorithms were proposed to attack general language classifiers. We adapt these five attacks by replacing their goal functions with Equation (1). These 25 attack recipes cover a wide range of popular transformations, constraints, and search methods (details in Table A1).
@@ 3.1 LLM Judge and Small Models based Automated Quality Controls
Removing text of low naturalness. Next we remove samples with low text naturalness using an English acceptability classifier (Proskurina et al., 2023). This classifier is fine-tuned from Huggingface TDA-BERT using a $3k$ labeled data we collect through human annotation. The human annotation guidelines on what defines "text naturalness" are in Section 3.2. We fine-tune the model with $2,370$ labeled texts, and evaluate it with $593$ held-out texts, following training setup in Section A.3. Table A3 shows that the F1 score ($88.9\%$) of fine-tuned TDA-BERT improves 18% compared to F1 ($70.5\%$) from pretrained TDA-BERT.
LLM judge for Removing non-toxic invalid TAE samples. Now we design model-based automated strategy to keep only those TAE samples that are semantically toxic. We propose an ensemble approach for toxicity label filtering by combining : (1) in-context learning (ICL) prompted Mistral (Mistral-7B-Instruct-v0.1) Jiang et al. (2023) and (2) a fine-tuned toxigen-RoBERTa classifier Hartvigsen et al. (2022) (via "AND"). For (1), Mistral ICL, we run a series of experiments to select the best ICL prompt formatting according to He et al. (2024) and build 5-shot ICL prompting by selecting demonstrations from our TAE dataset (see the prompt in Table A5). The accuracy of best Mistral ICL prompting is 76%. For (2), we fine-tune Toxigen-Roberta with $3.2k$ human annotated data (see annotation guideline in Section A.2 and training set up in Section A.3) and achieve a F1 score of 94% (Table A4).
@@ 3.2 Human Evaluation to Annotate TAE on Toxicity and Naturalness
We use human annotators to curate the toxicity and text naturalness of subsets of generated TAE examples. Three human annotators are asked to review the toxicity and three annotators are asked to annotate the text naturalness. The final label is assigned by unanimous vote, where a fourth adjudicator resolves any disagreements. (1) Toxicity is defined as "issues that are offensive or detrimental, including hate speech, harassment, graphic violence, child exploitation, sexually explicit material, threats, propaganda, and other content that may cause psychological distress or promote harmful behaviors." (2) Text naturalness is defined as "text that could be plausibly written by a human even if it includes ‘internet language’ that is outside ‘school grammar’".
@@
Dataset Seeding Source Train Test Jigsaw - 1.48MM 185k Off-Tweet - 20k 2.5k Raw TAEs Jigsaw 529,880 271,805 OffensiveTweet 57,639 77,418 TaeBench Jigsaw 197,734 38,539 OffensiveTweet 12,857 15,989 TaeBench+ Jigsaw 199,244 40,114 OffensiveTweet 13,837 16,115
@@ 4.1 TAE Generation with Proxy Models and Seeding Datasets
Running TAE attacks needs a set of text inputs that are toxic as seeds (denoted as ${\mathbf{x}}$ in Equation (1) of Section 2.1). We use the following two datasets as seeds for our TAE attacks.
@@
A dataset derived from the Wikipedia Talk Page dataset [^3] Toxic Comment Classification Challenge, https://www.kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge. Wikipedia Talk Page allows users to comment, and the comments are labeled with toxicity levels. Comments that are not assigned any of the six toxicity labels are categorized as "non toxic". We can use this data for both binary and multi-label toxicity detection tasks.
@@
Davidson et al. (2017) use a crowd-sourced hate speech lexicon from Hatebase.org to collect tweets containing hate speech keywords. Each sample is labeled as one of three classes: those containing hate speech, those containing only offensive language, and those containing neither. This data is for multi-class toxicity detection.
@@
One important property of adversarial attacks is the ability of the attack to transfer from the model used in its development to attacking other independent models. Transferability occurs because deep learning models often learn similar decision boundaries and features. Therefore, perturbations and noise patterns that fool one model are likely to also fool other models trained on the same or similar datasets. Motivated by adversarial transferability, we build three local text toxicity models as target proxies and run 25 different TAE attack recipes (see Section 2.1) against them to generate a large-scale pool of unfiltered TAE dataset (940k samples in total). Details of these proxy models are in Table A2 and Section A.4.
@@ 4.2 TaeBench: a Large Set of Quality Controlled TAE Samples
In Table 2, we pass 936,742 raw TAEs through the proposed quality filtering pipeline. We are able to select 264,672 examples (28.30% as of the original examples) as the filtered set, and we call it TaeBench. TaeBench is distributed as a toxic adversarial example dataset under a CC-BY-4.0 license, with metadata including generation recipe, transformations, constraints, seed sample/dataset/split.
To validate filtering quality, we conduct human annotations by randomly sampling 200 TAEs from each filtering step. In Table 2, human validation shows that, after filtering, the toxicity ratios are improved by 5.64% in the selected examples ($94.17\%$) compared to unfiltered examples ($88.53\%$). The text naturalness ratios are improved by 6.36%, from ($79.63\%$) in the unfiltered examples to ($85.99\%$) in the selected examples.
@@ 4.3 TaeBench+: Benign Seeds Derived Adversarial Examples
TAE are semantic-toxic samples that fool toxicity detection models into making benign predictions. Essentially they are false negative predictions (assuming "toxic" is the positive class). Related, it is also interesting to understand and search for those semantic-benign samples that fool a target model into making toxic predictions. These samples belong to false positive inputs. We call them "benign adversarial examples (BAE)".
To search for BAE, we design its goal function as:
${\mathcal{G}}({\mathcal{F}},{\mathbf{x}}^{\prime}):=\{{\mathcal{F}}({\mathbf{x}}^{\prime})\neq b;{\mathcal{F}}({\mathbf{x}})=b\}$ (2)
where $b$ denotes the benign class. Starting from benign seeds (${\mathcal{F}}({\mathbf{x}})=b$), we perturb ${\mathbf{x}}$ into ${\mathbf{x}}^{\prime}$ by pushing the prediction of ${\mathbf{x}}^{\prime}$ to not be benign anymore. We can reuse the TAE attack recipes by keeping their transformation, search and constraint components intact, and replace the goal function into the above Equation (2).
Empirically, we run the 25 BAE attacks, obtaining 102,667 raw BAE examples (searching for BAE seems harder than searching for TAE). Table A8 shows how we conduct automated filtering following the same workflow as obtaining TaeBench. Differently, in the label-toxicity filtering step, we keep those benign-labeled BAE samples. Finally, we add the filtered BAE examples to create TaeBench+, a new variation of the TaeBench dataset. We provide the additional benefits of TaeBench+ in Section 5.3.
@@ 5.1 Benefit I: Benchmark Toxicity Detectors via Transfer Attacks
To evaluate the efficacy of the filtered TAE examples, we conduct transfer attack experiments to benchmark four SOTA toxicity classifiers: detoxify (detoxify-unbiased) (Hanu and Unitary team, 2020), Llama Guard444meta-textgeneration-llama-guard-7b (Inan et al., 2023), OpenAI Moderation API555text-moderation-007 from https://platform.openai.com/docs/guides/moderation/overview, and Nemo Guardrails (with GPT-3.5-turbo) (Rebedea et al., 2023). Using TaeBench in transfer attacks can save resources and minimize the effort needed to generate TAE examples plus with data quality guarantees. Also the transfer attack set up is indeed a (major) real-world use case of using TAE. In this black-box transfer attack setup, TAE are constructed offline (like what we have done using many existing TAE attack recipes to attack local proxy models), then get them used to attack a target victim model.
@@
5.2 Benefit II: Improve Toxicity Detection w. Adversarial Training
We also showcase how vanilla adversarial training with TaeBench can help increase the adversarial robustness of a toxicity detector against unseen attacks. Here, adversarial training introduces the TAE adversarial data into the training of a DistilBERT or detoxify model together with the Jigsaw Binary train split (see Table 3 for more dataset details).
Table 5 reports the impacts of using TaeBench for adversarial training. We train DistilBERT/detoxify models with: (a) Jigsaw-train only (No TAE); (b) Jigsaw-train + extra unfiltered TAE (TAE-Unfiltered); and (c) Jigsaw-train + TaeBench. We sample the unfiltered TAE data such that TAE-Unfiltered has the same size as TaeBench to have a fair comparison on model performance by removing the impact of data set size. We observe that the model trained with Jigsaw-train + TaeBench achieves significantly lower ASR (14.58% and 23.25% FNR for DistilBERT and detoxify respectively), being more robust than no adversarial training (74.99% and 54.28% ASR/FNR) or random sampling augmentation (16.55% and 22.92% ASR/FNR). These augmentations minimally impact Jigsaw test set classification metrics (<2% F1/AUC change in Table 5). Training setups are described in Section A.3.
@@ 5.3 Variation: Adding TaeBench+
Table 5 also shows that when augmenting training data with TaeBench+, the model achieves the lowest ASR (FNR) of 12.66% and 22.80% on TaeBench-test for DistilBERT and detoxify respectively. We further oversample the benign adversarial examples in TaeBench+ during augmentation (balanced TaeBench+) to balance toxic and benign adversarial example sizes. This reduces the ASR (FPR) on (TaeBench+)-test-benign to 53.02% and 3.97%. Combining FPR and FNR, the model trained on balanced TaeBench+ achieves the highest balanced accuracy of 65.35% and 86.55% on the TaeBench+ test set.